Solve 5|x + 1| + 2 = 12
Smaller answer (nearest integer):
Larger answer (nearest integer):

Answer by oscargut(2103)

5|x + 1| + 2 = 12
5|x + 1| = 10
|x + 1| = 2
x+1 = -2 or x+1 = 2
x = -3 or x = 1
I can help you with all of your questions at: mthman@gmail.com
Thanks